### v1

# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

Abstract
To create culturally inclusive vision-language models (VLMs), the foremost requirement is developing a test benchmark that can diagnose the models’ ability to respond to questions reflecting cultural elements. This paper addresses the necessity for such benchmarks, noting that existing research has relied on human annotators’ manual efforts, which impedes diversity and efficiency. We propose a semi-automated pipeline for constructing cultural VLM benchmarks to enhance diversity and efficiency. This pipeline leverages human-VLM collaboration, where VLMs generate questions based on guidelines, human-annotated examples, and image-wise relevant knowledge, which are then reviewed by native speakers for quality and cultural relevance. The effectiveness of our adaptable pipeline is demonstrated through a specific application: creating a dataset tailored to Korean culture, dubbed K-Viscuit. The resulting benchmark features two types of questions: Type 1 questions measure visual recognition abilities, while Type 2 assess fine-grained visual reasoning skills. This ensures a thorough diagnosis of VLM models across various aspects. Our evaluation using K-Viscuit revealed that open-source models notably lag behind proprietary models in understanding Korean culture, highlighting areas for improvement. We provided diverse analyses of VLM performance across different cultural aspects. Besides, we explored the potential of incorporating external knowledge retrieval to enhance the generation process, suggesting future directions for improving cultural interpretation ability of VLMs. Our dataset and code will be made publicly available.
Evaluating Visual and Cultural Interpretation:
The
K-Viscuit Benchmark with Human-VLM Collaboration
Yujin Baek11footnotemark: 11     ChaeHun Parkthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang2     Jaegul Choo1 1 KAIST AI     2 KT Corporation {yujinbaek,ddehun,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo,dschang}@kt.com
1 Introduction
The utility of vision-language models (VLMs) in bridging visual and textual data is well-recognized, offering advancements in applications such as image captioning and optical character recognition. However, most VLMs are predominantly trained on data from Western languages and cultures (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts (Liu et al., 2021; Yin et al., 2023, 2021; Romero et al., 2024). Cultural nuances significantly influence the interpretation of visual and textual information, impacting the accuracy and relevance of VLM outputs. Consequently, building VLMs that understand diverse cultural backgrounds is essential for enhancing inclusivity and ensuring fair performance across different demographics.
To develop VLMs that effectively understand the cultural contexts of specific countries, establishing a wide-ranging test benchmark is a critical first step. Such a benchmark is essential for diagnosing the model’s ability to accurately interpret and respond to culturally specific inputs. Notably, previous research has endeavored to create VLM test benchmarks for particular countries (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024).
Figure 1: Pipeline Overview.
However, these approaches have primarily relied on manually crafted questions, which present two major limitations. First, their process is time-consuming and labor-intensive, increasing the cost of the research and making it difficult to extend and apply to the cultural contexts of other countries. Additionally, the diversity of questions can be limited due to cognitive fixation, where humans find it challenging to generate a wide range of varied questions after settling on an initial idea (Ramos, 2020). This lack of diversity can hinder the accurate diagnosis of the model’s ability to generalize across different types of culturally nuanced inputs.
To address these limitations, we propose a semi-automated pipeline for constructing cultural VLM benchmarks that enhance the efficiency and diversity of culture-relevant visual question and answer generation as illustrated in Fig. 1.
This pipeline incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers leveraging carefully crafted guidelines, a small set of human-annotated examples, and image-wise relevant knowledge. These recommendations are then assessed and approved by native speakers, ensuring high quality and cultural relevance.
In this work, we applied the proposed pipeline to create a dataset for Korean culture, named K-Viscuit (Korean Visual and Cultural Interpretation Test). Nevertheless, our pipeline can be adapted and applied to other cultures analogously. The resulting benchmark dataset created using our proposed pipeline has two notable features. Firstly, it has the ability to measure nuanced cultural understanding through two types of assessments: Type 1, which measures visual recognition capability, and Type 2, which measures visual reasoning capability. Secondly, the multiple-choice questions in the dataset include four options, with distractors generated following carefully designed instructions to ensure high similarity among the options. This high similarity between options prevents models from using shortcuts to determine the correct answer, ensuring a more robust and comprehensive evaluation of cultural understanding in VLMs. These features underscore the effectiveness and thoroughness of the resulting benchmark dataset in addressing cultural nuances.
Using our benchmark data, we evaluated the performance of various open-source and proprietary VLMs. Our experimental results indicate that open-source models significantly lag behind proprietary models in the context of Korean culture. This assessment highlights key improvement areas and sets development goals for open-source models, with proprietary models as the reference. Our contributions are summarized as follows:
  • We propose a pipeline for building benchmarks to diagnose VLMs’ understanding of various cultures. It enables efficient creation of culturally diverse datasets through human-VLM cooperation.
  • Using our pipeline, we constructed the K-Viscuit Korean Culture VQA benchmark.
  • We utilized K-Viscuit to diagnose and analyze the performance of existing open-source and proprietary VLMs.
2 Related Work
Recent research has made significant strides in developing benchmarks to assess the ability of AI models, particularly VLMs and Large Language Models (LLMs), to understand and respond appropriately to various cultural contexts. These efforts are crucial given that many existing models are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), which constrains their performance in non-Western settings. On the side of LLMs, Kim et al. (2024) introduced a benchmark dataset, CLIcK, specifically for assessing cultural and linguistic intelligence in Korean language models. This dataset includes QA pairs designed to test the cultural knowledge of Korean language models. Wibowo et al. (2023) created COPAL-ID, a dataset that captures Indonesian cultural nuances for text-based commonsense reasoning, highlighting the need for culturally aware benchmarks.
VLMs require consideration of both image and text inputs to effectively reflect cultural contexts, as cultural nuances can arise from either source. Liu et al. (2021) proposed the MaRVL dataset, a multilingual visually grounded reasoning dataset that includes five different languages and cultures to diagnose inherent language and cultural biases in pre-trained models. Their work highlights the need for models to seamlessly interpret and integrate visual and textual information across diverse cultures, paving the way for more inclusive and globally aware AI systems. Yin et al. (2021) built the GD-VCR dataset, which tests visual commonsense reasoning influenced by geographical and cultural contexts. Romero et al. (2024) presented the CVQA, a multilingual visual question-answering (VQA) benchmark that incorporates various languages and cultures, providing a comprehensive tool for evaluating VLMs in diverse cultural contexts.
Previous efforts to create cultural benchmarks for VLMs, however, have heavily relied on human annotators, which can limit both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024). The efficiency and diversity of dataset creation can be significantly enhanced by leveraging AI models, particularly when combined with human collaboration. This approach not only improves the speed and scope of dataset generation but also maintains high-quality standards. Liu et al. (2022) demonstrated this in their study on building a natural language inference dataset, where they combined the generative capabilities of language models with human review. Analogous to how LLMs are heavily utilized in the creation of datasets for language-only tasks (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024), leveraging the inherent knowledge of AI models in building multimodal cultural benchmarks can yield diverse and culturally nuanced examples. In this work, by harnessing AI models, we have significantly enhanced the diversity and richness of the dataset, ensuring it better reflects the wide array of cultural contexts we aim to represent.
3 Data Construction Pipeline
We present our human-AI collaborated dataset construction pipeline to diagnose the understanding of VLMs in a target cultural field. In this work, we focus on Korean culture as a target cultural field. First, we provide an overview (section 3.1) and details of the pipeline. Then, we present the analysis of the resulting dataset, dubbed K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
3.1 Pipeline Overview
The pipeline is designed to comprise a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct choice. Native Korean speakers are involved in the dataset construction, and a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and encourage an effective annotation process. The generated samples are designed to cover various aspects of a target culture derived from daily life and necessitate multimodal reasoning to interpret both the image and the question accurately. All text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects. The dataset construction consists of four stages: 1) concept selection, 2) image selection, 3) question and options annotation, and 4) human verification. Fig. 1 illustrates the overall pipeline. The details in each stage are as follows.
3.1.1 Concept Categorization
Inspired by recent studies on constructing multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge related to various concepts that Korean natives encounter in daily life. Each concept should not be exclusive to a specific culture but should have some degree of universality. To define the list of such concepts, we refer to semantic concepts provided in the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015), following Liu et al. (2021). Based on the concepts in the resource and discussions, our dataset is based on top of the following ten concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
3.1.2 Image Selection
Based on the list of chosen concepts above, Korean native annotators manually collected images from the web. To avoid an excessive number of images containing specific objects, each object is limited to appearing a maximum of two times within any single category. This encourages a broad representation of concepts related to Korean culture. Moreover, following Liu et al. (2021), only images depicting concepts that could physically exist in everyday life were selected. Annotators are encouraged to creatively choose diverse and suitable images by using various web resources. Wikimedia Commons [^1] https://commons.wikimedia.org/wiki are used as the primary sources, and only CC-licensed images are selected.
3.1.3 Question Generation
Question Type
Based on the selected images, we annotate questions in the form of multiple-choice QA. To comprehensively evaluate the understanding of Korean culture, we categorize the questions into two types: visual recognition (Type 1) and reasoning (Type 2). The visual recognition questions ask about basic visual information (e.g., counting, color recognition) or the names of objects appearing in the image. The reasoning questions go beyond simple visual recognition and ask for fine-grained knowledge or reasoning processes related to the given image. For each image, a single Type 1 question and between one and four Type 2 questions were created. This categorization offers two main advantages. First, the Type 1 questions allow for a multifaceted assessment of how well a model understands an image and its embedded concepts. Second, the Type 2 questions ensure that our dataset comprehensively evaluates the understanding of various aspects of Korean culture, rather than merely falling back to simple object recognition.
# of samples 657
- Type 1/ Type 2 237/420
# of unique images 237
# of options 2628
# of unique options 2129
Avg. question length 13.5
- Type 1/ Type 2 10.1/15.5
Avg. option length 1.7
Table 1: Dataset statistics. The length of questions and options denotes the number of words.
Figure 2: Concept distribution of our dataset.
AI-assisted Question Annotation
We create questions and their options (including one correct answer and three distractors) by leveraging a powerful proprietary large VLM (i.e., GPT-4-Turbo). For each concept category, human annotators first write questions and options for at least three images in each category. These manually annotated examples serve as demonstrations for the VLM to generate questions and options along with additional information.
Figure 3: Dataset Examples. We present an image and two questions of different types for each concept category.
In detail, the VLM receives 1) the target image, 2) a set of human-annotated demonstration examples for other images, 3) a detailed guideline for annotation, and 4) an image-wise textual knowledge description. We include a description or general knowledge associated with the given image to enhance the diversity and relevance of the questions. This ensured that the VLM could generate questions grounded in real-world knowledge. One important note is that the guidelines emphasize promoting similarity among all four options for multiple-choice questions, a principle kept in mind while creating the human-annotated examples, inspired by Wang et al. (2023). All textual information is delivered to the VLM through a natural language prompt, and different annotation guidelines are used for generating questions in different types (i.e., visual recognition and reasoning). The text prompts for VLM are presented in Appendix A.
3.1.4 Human Verification
During our preliminary study, we observed that the model often generated inappropriate data samples. For instance, some questions are open-ended or do not induce a single clear answer. In other cases, generated questions are subjective or not sufficiently related to the given image or Korean culture. To ensure the quality of automatically generated datasets, human verification is performed on the generated samples from the VLM. The VLM is guided to generate five question-option sets per image. If none of the five generated samples for a single image meet the criteria, we ask the VLM to regenerate questions and answers. The generated samples deemed appropriate are included in the final dataset with minimal revisions if required. We note that this AI-assisted annotation approach offered several advantages: It significantly accelerated the annotation process and helped minimize unintended human biases. By including human verification on generated questions, we ensure both accuracy and cultural relevance.
3.2 K-Viscuit
Statistics
The detailed statistics of our final benchmark dataset are shown in Table 1. Our dataset comprises a total of 657 examples, with 237 Type 1 questions and 420 Type 2 questions, drawn from 237 unique images from 10 concept categories. The average word count for questions in Type 1 and Type 2 are 10.11 and 15.46, respectively. The average word count for all questions is 13.53 words. Each question is accompanied by four options, resulting in a total of 2,628 options with an average word count of 1.74 words. Fig. 2 illustrates the distribution of selected concepts of samples in our dataset.
Required Knowledge to Solve Questions
To better understand the characteristics of our dataset, we categorize the types of required knowledge to solve questions. We analyzed Type 2 questions using the GPT-4 model following Tong et al. (2024). All Type 2 QA pairs were provided, and the knowledge required to answer each question was summarized. As shown in Fig. 4, diverse cultural elements were considered. Detailed instructions for the categorization can be found in the Appendix B.1.
Figure 4: Required Cultural Knowledge.
Qualitative Examples
Fig. 3 presents sampled images, questions, and options along with their concept category and question type.
4 Experiments
We conduct experiments to evaluate various VLMs on our constructed dataset.
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19
Table 2: VLMs Evaluation Results. Celeb., Arch., and Agri. denote Celebration, Architecture, and Agriculture. The highest and the second highest scores in each column are highlighted in bold and underlined.
4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), and Idefics2-8B Laurençon et al. (2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), and Gemini-1.5-Pro (Reid et al., 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
Model Type 1 Type 2 Total
InstructBLIP-7B 45.57 53.81 50.84
InstructBLIP-13B 51.48 57.86 55.56
mPLUG-Owl2-7B 43.04 51.19 48.25
LLaVA-1.6-7B 50.21 59.76 56.32
LLaVA-1.6-13B 54.01 58.81 57.08
InternLM-XC2-7B 56.12 61.67 59.67
Idefics2-8B 63.71 63.57 63.62
Claude-3-opus 69.62 80.24 70.02
GPT-4-Turbo 78.90 81.90 80.82
Gemini-1.5-Pro 83.97 70.24 81.58
Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and visual reasoning, respectively.
4.2 Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The Gemini-1.5-Pro and GPT-4-Turbo achieve the highest and second-highest scores, respectively. Among the open-sourced models, Idefics2 and InternLM-XComposer2-VL usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions, except for Gemini-1.5-Pro and Idefics2-8B. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. Our dataset enables the evaluation and identification of such aspects where models can be improved.
Figure 5: Qualitative examples of selected VLMs (i.e., LLaVA-1.6-13B, InternLM-XC2-7B, GPT-4-Turbo, and Gemini-1.5-Pro) on sampled questions. The answer option is highlighted in the underline. The correct and incorrect choice of models are highlighted in green and red.
4.3 Analyses
Asking the VLM in Korean
In this work, we primarily focused on evaluating the VLM’s understanding of Korean culture and intentionally excluded their understanding of the Korean language. Thus, we designed the dataset to focus on the VLM’s multiculturality, independent of its multilingualism. However, cultural information about certain groups often exists in the language that persons in the group frequently use. In other words, VLMs trained in multilingual corpora might have learned about Korean culture through texts written in Korean. Therefore, we probe whether asking questions in Korean could improve the performance of VLMs on our dataset. To translate our dataset into Korean, we use proprietary unimodal LM (gpt-3.5-turbo) as a machine translation system. For each sample in our dataset, the LM translates questions and four options into Korean. Three proprietary VLMs that can receive Korean text (i.e., GPT-4-Turbo, Claude-3-opus, and Gemini-1.5-Pro) are used for experiments.
Evaluation results with different input languages are presented in Table 4. We observe that solely providing translated texts in Korean to VLMs does not contribute to model performance. When English texts are also given to models, Gemini-1.5-Pro shows increased performance (i.e., 81.58 to 83.41). GPT-4-Turbo and Claude-3-opus usually do not take the benefits of using Korean texts.
Model en ko en + ko
Claude-3-Opus 70.02 65.30 70.32
GPT-4-Turbo 80.82 75.34 80.37
Gemini-1.5-Pro 81.58 80.82 83.41
Table 4: Results with Different Input Languages. ko is machine-translated texts. For en+ko, we provide questions and options in both languages to models.
Human Evaluation on the Benchmark
To evaluate how well people from different backgrounds perform on K-Viscuit, we conducted human evaluation. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 Non-Koreans. As depicted in Fig. 6, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist both between individuals and within a single person’s knowledge across different domains, such as history. Proprietary VLMs can match or exceed human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix B.2.
Figure 6: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on a K-Viscuit subset, measured across different proficiency levels.
Qualitative Results
Fig. 5 presents the prediction of selected models on sampled examples. In the first example, all models fail to correctly answer the question about asking detailed game rules. In the second case, two proprietary models are wrong while open-source models make correct answers. The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a similar function to the Griffith Observatory in the United States. Both proprietary models make correct answers to this example. In the final example, all models successfully identify the correct answer about agriculture.
Figure 7: Qualitative results of a VLM with multi-choice and generative VQA setups. LLaVA-1.6-13B is used for the analysis. The correct and incorrect model generations are manually highlighted by the authors.
5 Discussion
We discuss several components and future directions to build VLMs grounded on diverse cultures.
5.1 Evaluation beyond Multiple-choice VQA
Our dataset is designed as a multiple-choice VQA task, where VLMs are required to choose one of the given options. While this classification setup is advantageous for measuring model performance through accuracy, it may not fully capture the models’ understanding of Korean culture. For instance, VLMs may correctly select an answer from the given options but fail in a generative VQA setup where they must generate a free-form answer. An example of this is shown in Fig. 7 (left), where the VLM correctly identifies an object from the choices but hallucinates it as “stones or bricks” when generating an answer without options. In contrast, the right example in Fig. 7 shows the VLM accurately identifying the object as “bibimbap” in both multiple-choice and generative setups. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
5.2 Augmenting VLMs with External Knowledge
In previous experiments, we observed significant performance gaps in models for questions that require cultural knowledge. To address these gaps, two methodologies can be considered: fine-tuning open-source models and leveraging external knowledge. While fine-tuning is limited to open-source models, leveraging external knowledge can be applied to both open-source and proprietary models. Therefore, we focused on the latter approach.
We augmented the models with relevant documents for each test image from the Food concept. Our retrieval method involved generating captions using GPT-4-Turbo, embedding these captions via text-embedding-3-large to build queries, and retrieving Top-1 document from 152 Wikipedia pages related to the objects mentioned in the Type 1 options. The distractors in K-Viscuit were designed to be highly similar to the correct answers. This creates a challenging retrieval setting with numerous hard negatives, aiming to mimic real-world conditions. We aim to assess whether the retrieval method remains effective even in such a setting by observing the performance changes when cultural knowledge is added. The results show that providing retrieved documents can improve model performance, as indicated by higher scores in the Retrieved setting compared to None. The Oracle setting, which uses a curated relevant document, serves as an upper bound. Overall, augmenting VLMs with external knowledge via document retrieval demonstrates the potential to bridge the performance gap in VQA tasks requiring cultural context. We provide the details in the Appendix D.
Model None Retrieved Oracle
LLaVA-1.6-7B 43.66 68.31 78.87
LLaVA-1.6-13B 45.77 64.08 80.28
InternLM-XC2-7B 50.70 68.31 82.39
Idefics2-8B 51.41 67.61 82.39
Claude-3-opus 62.68 70.42 87.32
GPT-4-Turbo 73.94 78.17 88.73
Gemini-1.5-Pro 80.28 78.17 90.85
Table 5: Retrieval-augmented Generation Results.
6 Conclusion
In this work, we proposed a semi-automated pipeline to construct culturally aware benchmarks for vision-language models, featuring multi-choice QA with two types of questions that require visual recognition or reasoning. This pipeline combines human-VLM collaboration, enhancing efficiency and diversity in generating culturally relevant QAs. Our approach, first used to create the K-Viscuit dataset, can be adapted to other cultures.
Our evaluation with K-Viscuit identified a notable performance gap between open-source and proprietary VLMs in understanding Korean culture, highlighting the need for improvement in open-source models. Further analyses demonstrated the impact of in-depth cultural understanding on VQA results. Finally, we enhanced VLM performance through knowledge augmentation and explored open-ended answer generation beyond the multiple-choice setting. Our work stresses the importance of cultural diversity in evaluation, aiming for more inclusive and accurate VLMs in the future.
Limitations
Our current framework requires a manual selection of images that match specified concepts, preventing fully automated dataset generation. These human efforts can be alleviated by multimodal retrieval modules to some extent. To this end, it should be predetermined whether current multimodal encoders can sufficiently understand culturally nuanced images and texts. Enhancing retrieval models to better understand and match cultural contexts remains our exciting future work. The manual verification of automatically generated questions also can be a considerable burden. Developing a quality estimation module for generated questions could assist in this process by reducing the workload on human annotators.
Appendix A Dataset Construction Details
Guidelines to GPT-4-Turbo for Dataset Annotation
The detailed prompts for the annotation with GPT-4-Turbo are presented in Table 6 and Table 7.
Prompt for Type 1 annotations:
[System Prompt]
You are a helpful Korean annotator to make visual question answering datasets.
[User Prompt]
Given an image, generate a question asking for the name of the main object or the main activity that people are engaged in, and generate one correct option (answer) and three wrong options (distractors).
Detailed guidelines are as follows: 1. The objects shown in the image is called ‘‘{object_name}’’ in Korean. You can include this word into your correct options after translation into English. 2. All options should be written in up to 5 words. 3. Please struggle to make creative or challenging distractors so that they are not easily distinguished from the answer. 4. Distractors should seem similar to the correct answer and related to the category of the main object in image (e.g., Hanok - Agungi, Sarangchae, Anchae, Daecheongmaru). Distractors are better when they have similar color, shape, or texture with the answer. 5. Separate each distractor with “;” symbol. 6. Don’t make any explanation. 7. Distractors should be culturally related to the image. 8. All the options should be either transliterated or translated. Never mix transliteration and translation. 9. Don’t be too specific (Avoid using a proper name: instead use [University building] instead of [Ewha Campus Complex]) You can refer to below examples that are annotated for other images.
Question: What is the name of this place shown in the image?
Answer: Sarangchae Distractors: Anchae ; Sadang ; Daecheongmaru Question: What is the name of the structure seen in the image? Answer: Ondol Distractors: Agungi ; Jangdokdae ; Buttumak Question: What is the name of this building shown in the image? Answer: Gosiwon Distractors: Officetel ; Apartment ; Share house Please make four options (single answer and three distractors).
Table 6: A prompt of GPT-4-Turbo used in the annotation of Type 1 questions (i.e., visual recognition) in the Architecture category.
Prompt for Type 2 annotations:
[System Prompt]
You are a helpful Korean annotator to make visual question answering datasets.
[User Prompt]
Please ask 5 questions and their options about the image. Here are the guidelines to follow for writing.
Detailed guidelines are as follows:
1. The objects shown in the image is ‘‘{object_name}’’. Don’t include this word in your questions.
[…]
You can refer to below examples that are annotated for other images. Question: Seen in the image, what traditional Korean heated floor system is associated with the heat source from this feature? Answer: Ondol Distractors: Daecheongmaru ; Anchae ; Buttumak Question: In the image, what kind of Korean roof finishing is visible, known for its multicolored patterns? Answer: Dancheong Distractors: Seoggarae ; Cheoma ; Maru Question: Which mode of transportation is commonly used by tourists to ascend the mountain where the tower is located? Answer: Cable Car Distractors: Bus ; Bicycle ; Funicular Railway Please make four options (single answer and three distractors).
Table 7: A prompt of GPT-4-Turbo used in the annotation of Type 2 questions (i.e., visual reasoning) in the Architecture category.
Appendix B Dataset Analyses Details
B.1 Required Knowledge Analysis
We analyzed how diverse the cultural knowledge required by the questions in our K-Viscuit dataset is. To this end, we used the following prompt to obtain responses from the GPT-4 model. We delivered all Type 2 samples (including both the questions and the options) to the model by concatenating them into a single string.
Prompt for required knowledge analysis:
I created a multiple-choice quiz with four options per question, based on images related to Korean culture. Each question is designed to assess the understanding of one or more cultural elements. Please analyze which cultural element each question aims to measure and provide an overall summary. ‘‘{TYPE 2 samples}’’
Table 8: Prompt for required knowledge analysis.
B.2 Human Evaluation Details
We randomly selected images according to the proportion of each category to create the questionnaire for the human evaluation. If there were multiple Type 2 questions for a single image, we sampled them randomly. The number of selected images per category is as follows: Food (4), Beverage (2), Game (2), Celebrations (2), Religion (2), Tool (3), Clothes (2), Heritage (2), Architecture (4), and Agriculture (2).
We released the survey on the Amazon MTurk platform, where non-Koreans with a relatively limited understanding of Korean culture were asked to complete the K-Viscuit questions within 20 minutes for a compensation of \$5. The survey on the MTurk platform resulted in a demographic composed entirely of Americans. Their self-assessed proficiency levels were: Very familiar (35.7%), Somewhat familiar (50%), Slightly familiar (14.3%), and Not familiar at all (0%). For Koreans, we administered the survey to 20 graduate students in their mid-to-late twenties. We received feedback that Koreans had the most difficulty with questions related to history.
Appendix C VLM Evaluation Details
Model Implementation Details
We present further implementation details in VLMs used in our experiments. All open-source VLMs are implemented with the Transformers framework (Wolf et al., 2020), and the checkpoints are downloaded from Huggingface Hub [^2] https://huggingface.co/models. For proprietary models, gpt-4-turbo-2024-04-09, gemini-1.5-pro, and claude-3-opus -20240229 are used. The text prompt used for proprietary models is presented in Table 9.
Inference prompt for proprietary VLMs:
[System Prompt]
You will be given an image taken in Korea and a 4-way multiple-choice question. Answer the question based on the given image and your knowledge about Korean culture.
[User Prompt]
Question: ‘‘{question}’’
Options:
a. ‘‘{option_a}’’
b. ‘‘{option_b}’’
c. ‘‘{option_c}’’
d. ‘‘{option_d}’’
Make sure to respond with the option’s letter: ‘a.’, ‘b.’, ‘c.’, or ‘d.’. Do not make any additional explanation.
Table 9: Inference prompt for proprietary VLMs.
Appendix D Retrieval Methodology Details
For external knowledge retrieval, we generated image captions using GPT-4 with the prompt shown in Table 10.
Prompt for generating image captions:
You are an AI language model specializing in identifying and describing Korean food from photographs. When given a photograph of Korean food, your task is to accurately describe the food based on its visual characteristics and visible ingredients. Your description should include the name of the dish, main ingredients, common accompaniments, and notable features that help identify the food. Be detailed yet concise, providing clear and helpful information to those trying to understand Korean cuisine. Ensure that your description is within 150 words.
Table 10: Prompt for generating image captions.